OpenStudy (anonymous):

What is the value of z in the equation 4(2z + 3) = 12? −13 12 0 12

pooja195 (pooja195):

Distribute 4(2z + 3) = 1 4*2=? 4*3=?

OpenStudy (anonymous):

12

pooja195 (pooja195):

8z+12=12 we subtact 12 from both sides 8z=0 whats 0/8 ?

OpenStudy (anonymous):

0

pooja195 (pooja195):

Good therefore z=0